OpenStudy (anonymous):

The density of an object with a volume of 60.0 mL and a mass of 400.0 is ____ ?

OpenStudy (anonymous):

The answers are: A 0.150 g/mL B. 2.89 g/mL C. 5.50 g/mL D. 6.67 g/Ml

jimthompson5910 (jim_thompson5910):

density = mass/volume D = M/V D = ???

OpenStudy (anonymous):

we know that,D = M/V= 400/60=6.67(ANS)

OpenStudy (anonymous):

Lol yeah I forgot to close i figured it out earlier but thankssss